OpenStudy (anonymous):

if y=x^2+5x-9 what will the y value be if x = -2 is it -23 or -15 ? so i guess is -2^2 different from (-2)^2

OpenStudy (anonymous):

\[ (-2)^2 = (-1)^2(2)^2 = 1 * 4 = 4\] :)

OpenStudy (anonymous):

\[ \textrm{but remember that} \ (-1)^n \ \textrm{is only positive when n is an even number} \\ \ \\ (-1)^1 = -1 \\ (-1)^2 = 1 \\ (-1)^3 = -1 \\ (-1)^4 = 1 \\ \textrm{etc...}\]

OpenStudy (anonymous):

so if \[y= x^2+5x+6 \] and -4>= x >=0 then (-4,2) (-3,0)(-2,0)(-1,2)(0,6) right ?

OpenStudy (anonymous):

Your domain doesn't really make sense... but all of those points are on that parabola :)

OpenStudy (anonymous):

did you mean-4 ≥x ≤ 0 ?

OpenStudy (anonymous):

yes, thanks. I need to watch that !!!!

OpenStudy (anonymous):

Thanks! And the ≤ and ≥ are always a pain :P

OpenStudy (anonymous):

yes, not only that but even when i get an equation out that looks like it was designed by a rocket scientist, I plough through and impress myself except if I have to take a positive number to the other side, i make it stay positive....every time!!!! nearly :)

OpenStudy (anonymous):

then you have to go back like 50,000 steps trying to figure out where it happened? Yah, been there >.> At least you recognize it though! :D

OpenStudy (anonymous):

And kudos on plowing through rocket science! ^^

OpenStudy (anonymous):

haaa.. yes, its tough when you look at an equation and you have to get rid of squares,roots brackets negative powers and stuff flying every which way but I see a +3 in there somewhere, and I say to myself, I'm defiantly going to screw this up !!! haaa :) thanks for helping with the question. I just liked the look of the graph that was wrong better so I became unsure :)

OpenStudy (anonymous):

Anytime ^^